Starting from rest, an object accelerates on a straight level surface at the rate of 2 m/s2 for 10 s.What is the speed of the object at the end of 10 s?


1
Expert's answer
2022-01-10T09:16:39-0500
"v=v_0+at=0+2\ \dfrac{m}{s^2}\times10\ s=20\ \dfrac{m}{s}."